OpenStudy (anonymous):

Use the x-intercept method to find all real solutions of the equation. 14x^3-53x^2+41x-4=-4x^3+x^2+x+4

OpenStudy (jmark):

Given 14x^3-53x^2+41x-4=-4x^3+x^2+x+4 18x^3-54x^2+40x-8=0 let (x-2) is one of the factor use long division of (x-2) into 18x^3-54x^2+40x-8 left with 9x^2-9x+2 9x^2-6x-3x+2 3x(3x-2)-1(3x-2) (3x-2)(3x-1) hence x=2 x=1/3 x=2/3 ICSE 9th maths question papers - http://icse.edurite.com/icse-question-papers/icse-maths-question-papers-for-class-9.html